Calculate ΔG at 298 K if the partial pressures of NO2 and N2O4 are 0.40 atm and 1.64 atm , respectively. please explain. thanks in advance
Answer #1

2NO2(g) 040 atm → N204 (8) 1:64 atur AG= ? Acº (298K) = -5.40 kJ = -5400 J . KB = 16+ (0-4)2 Kf = . 10-25 TAG= Ago + RT lnky